# Fragment-Masked Molecular Optimization

### v2

@@ Introduction
For time $t$, we introduced objectives (Song et al. 2020) to generalize score matching and estimate the scores as follows:
@@
where, $\lambda_{1}(t)$ and $\lambda_{2}(t)$ are positive weighting functions and $B_{\phi}$ and $B_{\theta}$ denoted the noise prediction models based on the GNNs (Jo, Lee, and Hwang 2022; Zhang, Rao, and Agrawala 2023) to estimate scores $\nabla_{A}\mathrm{log}p_{t}(X_{t},A_{t},\mathbf{c})$ and $\nabla_{X}\mathrm{log}p_{t}(X_{t},A_{t},\mathbf{c})$, respectively. These two noise prediction models are jointly referred to as $\epsilon_{\theta}(\mathbf{G},\mathbf{c})$.
Fragment Mask Generation
Scaffolds typically refer to a molecule’s core structure or main ring system that determines its basic shape and properties (Li et al. 2019). In contrast, side chains are the branches or functional groups attached to the scaffolds. By altering the side chain properties, we can modulate the molecule’s solubility, polarity, reactivity, and various properties. As a result, we designed a fragment-based molecular optimization method. Molecular fragmentation is primarily based on the scaffolds and side chains in the molecule’s structure, and we applied rule-based constraints to the fragmentation results.
@@
In this instance, these side chains are referred to as fragments. After excluding the fragments containing only single atoms (e.g., C, N, Cl, and F), we verified the connectivity between $\mathcal{S}_{\mathbf{f}}$ and $\mathcal{S}_{\mathbf{c}}$ using $\mathcal{F}_{check}(\cdot)$.
@@
Fragment-Masked Molecular Optimization
@@
Specifically, as each reverse step from $\mathbf{G}_{t}$ to $\mathbf{G}_{t-1}$ relies solely on $\mathbf{G}_{t}$, it is essential to guide the masked region generation according to the known regions of $\mathbf{G}_{t}$ and the input optimization targets, described as follows:
@@
Table 1: Comparison of several methods for drug design across different cell lines for the PDD task. Success Rate represents the proportion of cell line types in which at least one optimized molecule is found.
Methods Cell: 909729 Cell: 687800 Cell: 687787 Total Cell (945)
N@100 Improv. N@100 Improv. N@100 Improv. Success Rate Improv.
GDSS (Jo, Lee, and Hwang 2022) - 0% - 0% - 0% 0.8% 1.1%
DiGress (Vignac et al. 2023) 3 3.4% 4 6.4% 1 4.2% 69.3% 3.0%
DruM2D (Xu et al. 2023) 1 4.3% 2 6.5% 2 4.2% 69.4% 2.6%
MOOD (Lee, Jo, and Hwang 2023) 1 0.7% 7 3.8% 1 0.6% 34.8% 1.7%
CDGS (Huang et al. 2023) 3 0.8% 3 3.9% 3 1.6% 32.5% 1.6%
FMOP (Ours) 11 7.8% 10 5.6% 15 8.8% 94.4% 5.3%
@@
Rule-Based Chemical Bond Post-Processing. During the sampling phase, the GNN-based score prediction model’s inherent limitations in the molecular generation process prevent each atom from obtaining information beyond the GNN layer’s $k$-hop neighborhood. This limitation may cause the model to miss global features or long-range interactions between atoms when generating molecular structures, which can, in turn, affect the overall structural rationality and stability of the molecule (Kim et al. 2023; Cai et al. 2022). Therefore, global optimization is necessary.
Thus, we changed the global features of the sampled molecule $\mathbf{G}_{0}$ to correct specific local edge types. Specifically, we changed the molecular structure by converting continuous double bonds to single and modifying six-carbon double-bonded to aromatic rings. These corrections ensure that the generated molecules meet the chemical stability and structural rationality requirements, thereby improving the accuracy and effectiveness of molecular optimization.
Experiment
Figure 3: Visual comparison of our optimization method with generative methods. This illustrates the unique molecular structures generated by our optimization method and compares them with various baselines across five distinct cell lines. Our method consistently produces diverse and effective molecules tailored to each cell line, avoiding convergence to the same local optimum.
Experimental Setup
@@
This study utilized two primary datasets: QM9 (Ramakrishnan et al. 2014) and GDSCv2 (Yang et al. 2013). The QM9 dataset was used for pre-training the model to enhance molecular diversity, and contains approximately 133,885 molecules. These data provide the model with rich molecular information, improving its generalization ability. Likewise, the GDSCv2 dataset was used for tasks related to drug response prediction (Shubham et al. 2024; Campana, Prasse, and Scheffer 2024), and comprises approximately 190,853 samples, covering 945 cell lines and 220 drugs. GDSCv2 enables learning and predicting drug response distributions. This is significant for precision medicine and new drug development (Ren et al. 2023).
@@
Evaluation Criteria. Molecules labeled with drug response for the cell line $\eta$ are selected from the GDSCv2 dataset, focusing on those with IC50 values in the top 20% to 30%. These molecules, referred to as $y_{\eta}$, were used as the target molecules for optimization. N@100 is a counting function, if the IC50 of the optimized molecules $y_{\eta}^{{}^{\prime}}$ is lower than the average $y_{\eta}$, the count is incremented by one. The Improv. represented the improvement in IC50 before and after molecular optimization. Our method calculated the average increase based on the true IC50 values for each molecule being optimized, while other methods also used the average IC50 of these molecules to compute the average increase.
@@
It is important to note that the generated or optimized molecules’ the efficacy (i.e., the IC50) in various cell lines was predicted using the deep learning models and not wet lab validation. This is because the wet lab validation involves molecular synthesis and cell-based assays, processes that are time-consuming and extremely costly. Therefore, we relied on the out-of-domain drug response prediction (OOD-DRP) methods (Li et al. 2024c; Chen et al. 2022; Sharifi-Noghabi et al. 2021) to predict the IC50 of the newly generated molecules. The contrastive learning drug response model based on natural language supervision (CLDR) focused on out-of-domain generalization and demonstrated state-of-the-art correlation in response predictions (Li et al. 2024b). Consequently, we used the CLDR as the OOD-DRP method.
Overall Experiments
To validate the effectiveness of FMOP, this study has meticulously ensured fairness by comparing it with five baseline models. GDSS (Jo, Lee, and Hwang 2022), DiGress (Vignac et al. 2023), and DruM2D (Xu et al. 2023) are unconditional molecular generation methods, while MOOD (Lee, Jo, and Hwang 2023) and CDGS (Huang et al. 2023) are conditional molecular generation methods that utilize specific category and the dataset distributions to guide molecular generation. Existing molecular optimization methods can be broadly categorized into target pocket- and property-based molecular optimization. Target pocket-based methods require detailed structural information and binding site data, whereas property-based optimization methods fail to integrate cellular genomics and drug-cell fusion information. Consequently, these methods cannot be directly compared with our approach.
To verify whether our method can effectively optimize molecules to achieve better drug response values, we conducted overall experiments involving various diffusion models. Due to the novelty of the molecule generation method, generated molecules are out-of-domain and require the OOD-DRP model to have a high generalization capability. Therefore, we used the CLDR method (Li et al. 2024b), which has excellent generalization performance, to predict the optimized molecules.
Table 1 displays the optimization results of six methods in three different cell line scenarios and the average optimization results across 945 cell lines. Our method achieves the best optimization and increase rate results, surpassing the second-best method by 36.0% and 76.7%, respectively. To verify that our method can effectively optimize fragment regions while maintaining scaffold consistency, we conducted a visual comparison of the results. As shown in Fig. 3, our method maintains scaffold consistency while optimizing the masked region. In contrast, other methods’ optimization success was primarily due to the random generation of a few molecules with good efficacy across various cell lines. However, these methods cannot adjust the sampling space distribution for specific tasks, resulting in poor performance.
Ablation Study
@@
Table 2: The ablation study shows the impact of different components on success rate, improvement, and total N@100 (TN@100).
@@
Origin Method 94.4% 5.3% 23789
(a) For the cell line 906792
(b) For the cell line 688006
(c) For the cell line 949090
Figure 4: Visualizations results for the IC50 distribution of molecules generated by comparative methods.
@@
As shown in Table 2, each component significantly contributes to the model’s overall performance. For Q1, without the fragment mask prompt, our method generates molecules randomly, resulting in a drastic decline in the success rate to 0.3% and an improvement of 2.1%, with only 26 instances reaching N@100. This indicates that the fragment mask is crucial for identifying key molecular features.
For Q2, when task guidance is removed, the success rate is 5.0%, and the improvement rate remains at 2.1%. The absence of task guidance leads to a random fragment generation strategy, where we use the top $N$ most frequent fragments to assemble the scaffold, yielding a success rate of 5%. This demonstrates that task guidance is essential for potential molecule optimization.
@@
For Q3, the success rate further increased to 94.4% through post-processing based on the aromatic ring recognition rules. This suggests that molecular modifications are essential for achieving a high success rate, although they introduce some complexity.
Visualization Analysis

### v3

@@ III-A Molecular Conditional Generation
For time $t$, we introduced objectives [53] to generalize score matching and estimate the scores as follows:
@@
where, $\lambda_{1}(t)$ and $\lambda_{2}(t)$ are positive weighting functions and $B_{\phi}$ and $B_{\theta}$ denoted the noise prediction models based on the graph neural networks (GNNs) [35, 54] to estimate scores $\nabla_{A}\mathrm{log}p_{t}(X_{t},A_{t},\mathbf{c})$ and $\nabla_{X}\mathrm{log}p_{t}(X_{t},A_{t},\mathbf{c})$, respectively. These two noise prediction models are jointly referred to as $\epsilon_{\theta}(\mathbf{G},\mathbf{c})$.
III-B Fragment Mask Generation
Scaffolds typically refer to a molecule’s core structure or main ring system that determines its basic shape and properties [55]. In contrast, side chains are the branches or functional groups attached to the scaffolds. By altering the side chain properties, we can modulate the molecule’s solubility, polarity, reactivity, and various properties. As a result, we designed a fragment-based molecular optimization method. Molecular fragmentation is primarily based on the scaffolds and side chains in the molecule’s structure, and we applied rule-based constraints to the fragmentation results.
@@ III-B Fragment Mask Generation
In this instance, these side chains are referred to as fragments. After excluding the fragments containing only single atoms (e.g., ’C’, ’N’, ’Cl’, and ’F’), we verified the connectivity between $\mathcal{S}_{\mathbf{f}}$ and $\mathcal{S}_{\mathbf{c}}$ using $\mathcal{F}_{check}(\cdot)$.
@@
III-C Fragment-Masked Molecular Optimization
@@
Specifically, as each reverse step from $\mathbf{G}_{t}$ to $\mathbf{G}_{t-1}$ relies solely on $\mathbf{G}_{t}$, it is essential to guide the masked (unknown) region generation according to the known regions of $\mathbf{G}_{t}$ and the input optimization targets, described as follows:
@@
Rule-Based Chemical Bond Post-Processing. During the discretization of the sampled molecular graph, discretization errors may occur with continuous edge features, which could result in the generation of chemically unreasonable or unstable structures. Furthermore, the GNN-based score prediction model’s inherent limitations in the molecular generation process prevent each atom from obtaining information beyond the GNN layer’s $k$-hop neighborhood. This limitation may cause the model to miss global features or long-range interactions between atoms when generating molecular structures, which can, in turn, affect the overall structural rationality and stability of the molecule [56, 57]. Therefore, global optimization is necessary. To address this issue, we employed a rule-based chemical bond post-processing method. This method automatically detects and corrects potential structural inconsistencies in the generated molecules after the initial structure generation process.
TABLE II: Comparison of several methods for drug design across different cell lines for the PDD task. Success Rate represents the proportion of cell line types in which at least one optimized molecule is found.
Methods Cell: 906792 Cell: 687800 Cell: 684055 Cell: 908149 Total Cell (985)
N@100 Improv. N@100 Improv. N@100 Improv. N@100 Improv. Success Rate Improv.
VAE-based JT-VAE [33] 18 2.89% 18 4.07% 9 5.69% 8 2.74% 92.70% 3.4%
Graph-based GeoLDM [34] 1 4.30% 2 6.50% 1 3.88% 1 3.46% 69.40% 2.6%
DiGress [41] 3 3.40% 4 6.40% 2 0.12% 3 2.76% 69.30% 3.0%
Diffusion-based GDSS [35] - - - - - - - - 0.80% 1.1%
MOOD [40] 1 0.70% 7 3.80% 1 0.51% - - 34.80% 1.7%
CDGS [39] 3 0.80% 3 3.90% 1 0.01% 3 0.61% 32.50% 1.6%
Fragment-based MARS [13] 4 5.63% 2 10.25% 4 5.76% 19 6.02% 93.50% 6.1%
FFLOM [38] 12 7.06% 6 12.97% 1 6.62% 16 8.61% 88.83% 6.3%
DST [12] 44 2.63% 22 4.97% 7 5.16% - - 55.53% 4.1%
Prompt-MolOpt [10] 9 7.42% 2 1.36% 2 1.36% 3 6.88% 91.68% 5.7%
HN-GFN [32] 12 3.55% 10 5.99% 8 5.74% 14 3.74% 92.70% 3.3%
FMOP (Ours) 15 7.80% 62 5.57% 67 4.58% 85 9.53% 95.43% 7.5%
(a) For the cell line 906792
(b) For the cell line 687800
(c) For the cell line 684055
(d) For the cell line 908149
Figure 3: Visualizations results for the IC50 distribution of molecules generated by fragment-based methods.
(a) For the cell line 906792
(b) For the cell line 687800
(c) For the cell line 684055
(d) For the cell line 908149
Figure 4: Visualizations results for the IC50 distribution of molecules generated by graph- and diffusion-based methods.
Specifically, we modified the molecular structure in the following ways:
  1. 1.
    Conversion of continuous double bonds to single bonds: This change is intended to prevent the formation of unstable chemical structures. In certain chemical reactions, consecutive double bonds can give rise to reactive intermediates that are both unstable and highly reactive.
  2. 2.
    Modification of six-carbon chains with double bonds into aromatic rings: This transformation improves stability, as aromatic rings are generally more stable than alkenes due to their conjugated electron structures.
Through this post-processing step, we not only improve the chemical stability of the generated molecules but also ensure their structural rationality. The rule-based method integrates chemical knowledge with computational models, resulting in molecules that better conform to actual chemical principles and drug development requirements, thereby enhancing the accuracy and effectiveness of molecular optimization.
IV Experiment
IV-A Experimental Setup
@@
This study utilized two primary datasets: QM9 [58] and GDSCv2 [19]. The QM9 dataset was used for pre-training the model to enhance molecular diversity, and contains approximately 133,885 molecules. These data provide the model with rich molecular information, improving its generalization ability. Likewise, the GDSCv2 dataset was used for tasks related to drug response prediction [59, 60], and comprises approximately 190,853 samples, covering 985 cell lines and 220 drugs. GDSCv2 enables learning and predicting drug response distributions. This is significant for precision medicine and new drug development [61].
@@
Evaluation Criteria. Molecules labeled with drug response for the cell line $\eta$ are selected from the GDSCv2 dataset, focusing on those with IC50 values in the top 20% to 30%. These molecules, referred to as $y_{\eta}$, were used as the target molecules for optimization. N@100 is a counting function, if the IC50 of the optimized molecules $y_{\eta}^{{}^{\prime}}$ is lower than the average $y_{\eta}$ and the reduction exceeds 1%, the count is incremented by one. The Improv. represented the improvement in IC50 before and after molecular optimization. Our method calculated the average increase based on the true IC50 values for each molecule being optimized, while other methods also used the average IC50 of these molecules to compute the average increase. Success Rate $=\frac{{\textstyle\sum_{i=1}^{\mathrm{M}}}\left\{\mathrm{N}_{i}@100>0\;\mathrm{and}\;\mathrm{Improv}(i)>1\%\right\}}{\mathrm{M}}$ represents the proportion of 985 cell line types (denoted as $\mathrm{M}$) in which at least one optimized molecule is found.
@@
It is important to note that the generated or optimized molecules’ the efficacy (i.e., the IC50) in various cell lines was predicted using the deep learning models and not wet lab validation. This is because the wet lab validation involves molecular synthesis and cell-based assays, processes that are time-consuming and extremely costly. Therefore, we relied on the out-of-domain drug response prediction (OOD-DRP) methods [23, 62, 63] to predict the IC50 of the newly generated molecules. When drug information is unavailable during the training phase and only cell line types are used for training, the testing accuracy typically ranges from the Pearson correlation coefficient [63] of 0.6 to 0.8. This accuracy depends on factors such as the distribution of the dataset, the diversity of cell lines, and the complexity of the predictive model used. The contrastive learning drug response model based on natural language supervision (CLDR) focused on out-of-domain generalization and demonstrated state-of-the-art correlation in zero-shot response predictions [22]. Consequently, we used the CLDR as the OOD-DRP method.
IV-B Baselines
To validate the effectiveness of FMOP, this study has meticulously ensured fairness by comparing it with eleven baseline methods:
  • JT-VAE [33]: A variational autoencoder (VAE) model [64] for molecular graph generation. It first generates a tree-structured scaffold library, and then combines selected samples from the scaffold library into the molecules using the graph message-passing network [65].
  • GeoLDM [34]: A latent diffusion model designed for 3D molecular generation, using autoencoders to encode structures into latent codes and diffusion models to operate in the latent space.
  • DiGress [41]: A discrete denoising diffusion model that iteratively adds or removes edges and modifies categories. A graph transformer network is trained to reverse this process.
  • GDSS [35]: A graph generative model based on score-based diffusion, utilizing a system of stochastic differential equations (SDEs) [66] to model the joint distribution of molecular nodes and edges. It generates molecules that adhere to chemical valency rules and closely follow the training distribution.
  • MOOD [40]: A score-based diffusion model for exploring chemical space, utilizing out-of-distribution (OOD) [67] control in the generative process to generate novel molecules. It is conditioned on target properties such as drug-likeness and synthesizability, guiding the diffusion process toward high-quality molecules.
  • CDGS [39]: A conditional diffusion model for molecular graph generation, incorporating OOD control in a generative SDEs to explore novel regions of chemical space.
  • MARS [13]: A multi-objective drug discovery method that iteratively edits molecular graph fragments using Markov Chain Monte Carlo (MCMC) [68] sampling and the GNNs.
  • FFLOM [38]: A flow-based autoregressive model for fragment-to-lead optimization, which generates molecular structures by linking fragments and growing R-groups.
  • DST [12]: A differentiable scaffolding tree method for molecular optimization, which converts discrete chemical structures into locally differentiable ones for gradient-based optimization.
  • Prompt-MolOpt [10]: A molecular optimization method that leverages prompt-based embeddings to enhance the transformer’s ability to optimize molecules for targeted properties.
  • HN-GFN [32]: A multi-objective Bayesian optimization (MOBO) [69] algorithm that uses a hypernetwork-based GFlowNets [70] (HN-GFN) as an acquisition function optimizer.
TABLE III: The table presents the optimized molecular properties of 12 molecular optimization methods, including molecular weight (MW, typically ranging from $>50$g/mol to $<500$g/mol), the log of the partition coefficient of a solute between octanol and water(LogP, typically ranging from $-1.5$ to $<5$), hydrogen bond donors (HBD, typically ranging from 00 to $5$), hydrogen bond acceptors (HBA, typically ranging from $1$ to $10$), polar surface area (PSA, typically ranging from $<20\ \text{Å}^{2}$ to $>140\ \text{Å}^{2}$), rotatable bonds (ROTB, typically ranging from 00 to $15$), and quantitative estimation of drug-likeness (QED, typically ranging from 00 to $1$).
      Methods       MW       LogP       HBD       HBA       PSA       ROTB       QED
      Base       452.25       3.6175       2.2332       6.4484       98.9591       5.7399       0.4641
      JVAE       305.47       2.5448       0.9769       3.7962       60.2302       4.5326       0.8134
      MOOD       283.39        5.6476        0.0469        0.3971        5.2514       4.3908       0.5848
      GDSS       128.17       0.2786       1.9167       2.0833       53.2942       3.1667       0.5702
      CDGS       122.20       0.7358       1.3031       2.0334       41.0165       2.8278       0.5386
      DiGress       123.47       0.4331       0.6862       2.6745       33.4768       1.1279       0.5241
      GeoLDM       113.80       0.3699       1.8906       2.0741       51.4864       1.0894       0.5103
      MARS       404.88       3.2676       1.7030       5.8044       84.4286       5.6015       0.4476
      Prompt-MolOpt        516.12       3.2727       2.6674       8.0261       126.2293       6.3037       0.3500
      DST       495.81       4.8649       2.4808       6.7039       101.4287       7.4716       0.3407
      FFLom        524.99       4.3199       1.8279       7.1002       99.6118       8.0676       0.3322
      HNGFN        533.64       3.4179       3.0587       8.7698        155.8988       5.0059       0.3322
      FMOP (Ours)       439.41       3.1507       3.2866       6.2079       96.8763       6.6725       0.3567
@@ IV-B Baselines
TABLE IV: The ablation study shows the impact of different components on success rate, improvement, and total N@100 (TN@100).
@@
Origin Method 95.4% 7.5% 23789
IV-C Overall Experiments
To verify whether our method can effectively optimize molecules to achieve better drug response values, we conducted overall experiments involving various diffusion models. Due to the novelty of the molecule generation method, generated molecules are out-of-domain and require the OOD-DRP model to have a high generalization capability. Therefore, we used the CLDR method [22], which has excellent generalization performance, to predict the optimized molecules.
It is crucial to highlight that the optimization task involving 985 cell lines is similar to optimizing across various properties. Our approach enables a single training process to address all tasks simultaneously, whereas other methods would require separate training for each of the 985 cell lines individually. Table II displays the optimization results of 12 methods in 4 different cell line scenarios and the average optimization results across 985 cell lines. Our method achieves the best optimization and increase rate results.
In order to confirm that our method can effectively optimize fragment regions while maintaining scaffold consistency, we conducted a visual comparison of the results. As shown in Fig. 5, our method maintains scaffold consistency while optimizing the masked region. Based on the distribution of cell line attribute features, FMOP employs regressor-free guidance, effectively generating molecules with specific attributes. The scaffold structures of these molecules are similar to the original ones being optimized, with the property values improved through side-chain optimization. In contrast, molecules generated by other optimization methods exhibit significant differences from the original molecules. Even fragment-based methods (DST, JT-VAE, and Prompt-MolOpt) still show poor performance. This is because these fragment-based methods essentially rely on statistical fragments from the training set to form a fragment library. During the molecular optimization phase, specific fragments are conditionally selected from the library and stitched together, but these methods cannot optimize fragments within the scaffold.
Some graph- and diffusion-based methods (GDSS, CDGS, DiGress, and GeoLDM) achieved success (1$\sim$3%) mainly by randomly generating a few molecules with good efficacy across various cell lines, which failed to adjust the sampling space distribution for specific tasks, resulting in suboptimal performance. To this end, we analyzed the molecular properties, such as MW, logP, QED, etc. In early drug design, the Lipinski’s rule of five [71], is commonly used as a guideline for evaluating drug candidates. We have marked results that fall outside the normal property ranges in red. As shown in Table III, the average MW and LogP of the molecules generated by these methods are significantly lower than those in the original dataset. Specifically, the original dataset has an average MW of approximately 452, while the molecules generated by these methods have an average MW of about 120. This indicates that the generated molecules have notably smaller molecular weights, which could potentially lead to adverse changes in their pharmaceutical properties.
IV-D Ablation Study
Figure 5: Visual comparison of our optimization method with generative methods. This illustrates the unique molecular structures generated by our method and compares them with various baselines across four distinct cell lines. Our method consistently produces diverse and effective molecules tailored to each cell line, avoiding convergence to the same local optimum.
@@ IV-D Ablation Study
Figure 7: Front-end visualization of the FMOP method on the Beishenglai platform. Users input the target conditions and submit the optimization task.
@@
As shown in Table IV, each component significantly contributes to the model’s overall performance. For Q1, without the fragment mask prompt, our method generates molecules randomly, resulting in a drastic decline in the success rate to 0.3% and an improvement of 2.1%, with only 26 instances reaching N@100. This indicates that the fragment mask is crucial for identifying key molecular features.
For Q2, when task guidance is removed, the success rate is 5.0%, and the improvement rate remains at 2.1%. The absence of task guidance leads to a random fragment generation strategy. In the QM9 dataset, there are 1798 fragments with a frequency of occurrence greater than 10, which we have collected into a fragment library. For each optimization task, we randomly select 100 fragments from this library. These fragments are then attached to the atom in the original molecule that has the most implicit hydrogen atoms. The atom types are limited to ’C’, ’N’, ’O’, ’S’, and ’P’. The first atom of each selected fragment is connected to the target atom via a single bond to prevent the covalent bond from exceeding the threshold set for the central atom, resulting in a success rate of 5.0%. This demonstrates that task guidance is essential for potential molecule optimization.
@@
For Q3, the success rate further increased to 95.4% through post-processing based on the aromatic ring recognition rules. This suggests that molecular modifications are essential for achieving a high success rate, although they introduce some complexity.
Overall, these findings highlight the importance of fragment masks, task guidance, and post-processing rules in enhancing the quality of molecular generation, particularly in improving the generation of aromatic rings.
IV-E Visualization Analysis